the ratio of the angle measures in a quadrilateral is 1:4:5:6. find each angle measure
Answer by checkley77(12844)

1:4:5:6
1+4+5+6=16
360/16=22.5
1*22.5=22.5 is the smallest angle.
4*22.5=90 is the next angle
5*22.5=112.5 is the third angle
6*22.5=135 is the largest angle
Proof:
22.5+90+112.5+135=360
360=360